OpenStudy (anonymous):

WILL GIVE MEDAL AND FAN!~~~A rock is dropped from a height of 2.5 meters. What is its velocity when it reaches the ground? 3.5 meters/second 7.0 meters/second 12 meters/second 14 meters/second 49 meters/second

OpenStudy (anonymous):

7 meter per second.

OpenStudy (anonymous):

its initial velocity is zero and its acceleration is g. u can now use the equation v^2-u^2=2as where v=final velocity u=initial velocity s=distance travelled a=acceleration

OpenStudy (anonymous):

\[2gh= V _{f}^{2} - V ^{2}_{i}\] here V f is the final velcoity which we have to find. and V i is the initail velocity. initail velocity is zero.

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

2(9.8) ( 2.5) = Vf ^ 2

OpenStudy (anonymous):

V ^2 = 49 By taking suqare root of both sides we have.... V = 7 m/s

OpenStudy (anonymous):

Thanks @Rohitkhanna @IhteshamMalik

OpenStudy (anonymous):

You're welcome